Circular No. 6411

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1915+105 S. N. Zhang, C. R. Robinson, B. A. Harmon, W. S. Paciesas, and G. J.

Fishman, Marshall Space Flight Center, NASA, report for the Compton Gamma Ray Observatory BATSE team: “A hard x-ray outburst from GRS 1915+105 is being detected by BATSE/CGRO. Its 20- to 100-keV flux has increased almost linearly from an upper limit of ∼ 50 mCrab on May 15–18 to ∼ 250 mCrab on May 24 and remained bright with significant day-to-day variations of 20–50 percent as of May 27. This flux level is slightly below the average peak flux (∼ 300 mCrab) of this source during previous outbursts.”

G. Pooley, Mullard Radio Astronomy Observatory, Cambridge, writes: “After four months of apparent inactivity in the radio band, GRS 1915+105 has shown a bright flare at 15 GHz. The Ryle Telescope has been monitoring the source intermittently since the outburst at the end of 1995 (IAUC 6269). The only significant detections since 1996 Jan. 20 were on the four days May 23–26, when peak flux densities near 100 mJy were observed. Strong quasiperiodic oscillations were observed, with periods near 40 min (cf. IAUC 6269) and 25 min, and possibly others. The source was only marginally detected on May 27 (4 mJy) and not detected on May 28.”

GRO J1655–40 R. M. Hjellming and M. P. Rupen report recurrence of radio emission

from the x-ray transient GRO J1655–40: “A Very Large Array observation at 4.8 GHz on May 29.25 UT showed the presence of a 19-mJy radio source at the position of the radio, optical, and x-ray source first detected in 1994. The radio source was not present (with upper limits of 0.5 mJy) in VLA observations from 1995 Jan. through 1996 May 20. This radio source is probably related to the recurrence of x-ray and optical activity reported on IAUC 6393 and 6406.”

SUPERNOVA 1996ae IN NGC 5775 R. M. Rich and R. A. Mendez report the following rough CCD magni-

tudes of SN 1996ae obtained on May 28.0 UT with the Cerro Tololo 0.9-m telescope: U = 18.7 ± 0.1, B = 18.8, V = 17.4, R = 18.3, I = 17.6 ± 0.05.

1996 May 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green